OpenStudy (pulsified333):

For this problem, assume 9 males audition, one of them being George, 6 females audition, one of them being Jackie, and 5 children audition. The casting director has 3 male roles available, 1 female role available, and 2 child roles available. How many different ways can these roles be filled if exactly one of George and Jackie gets a part?

OpenStudy (pulsified333):

@jim_thompson5910 @satellite73 @dan815

jimthompson5910 (jim_thompson5910):

`if exactly one of George and Jackie gets a part` so they mean just one of them? either George OR Jackie? The wording they have is a bit odd

OpenStudy (pulsified333):

yes, i know the wording is odd

jimthompson5910 (jim_thompson5910):

I'm going to assume they meant to say George OR Jackie (exactly one person) gets the role. So it's not possible for both to get the role

jimthompson5910 (jim_thompson5910):

`9 males audition` `3 male roles available` Let's say George is locked in to get a role. So we really have 3-1 = 2 slots left for the male roles. This is from a pool of 9-1 = 8 males how many ways are there to fill 2 slots from a pool of 8 males?

OpenStudy (pulsified333):

ok

OpenStudy (pulsified333):

56

jimthompson5910 (jim_thompson5910):

order doesn't matter. So you divide 56 by 2 to get 28

OpenStudy (pulsified333):

okay but isnt it possible to have any of the three roles

jimthompson5910 (jim_thompson5910):

I guess it depends on what the roles are. Are they important roles or extras? Hmm I'm not sure

jimthompson5910 (jim_thompson5910):

if the roles are important (eg: leading roles), then order would matter since the characters are different

OpenStudy (pulsified333):

so then how do we account for the three roles he could play

jimthompson5910 (jim_thompson5910):

I'm not sure now because I'm not sure what the roles are. So I don't know if order matters or not

jimthompson5910 (jim_thompson5910):

Sorry I'm not really helpful with this one

OpenStudy (pulsified333):

well i have 4 attempts so lets do it with three separate roles on this

jimthompson5910 (jim_thompson5910):

ok well if order did matter, then George could have male role A, male role B, male role C he has 3 choices if he picks male role A, then there are 8*7 = 56 ways to fill up the other 2 male roles if he picks male role B, then there are 8*7 = 56 ways to fill up the other 2 male roles if he picks male role C, then there are 8*7 = 56 ways to fill up the other 2 male roles as a shortcut, there are 3*56 = 168 ways to fill up the male roles assuming George gets a male role and order matters (the roles are important roles)

OpenStudy (pulsified333):

ok

jimthompson5910 (jim_thompson5910):

how many ways are there to fill the female role? We're still under the assumption George got the role. So Jackie can't get the role if George did.

OpenStudy (pulsified333):

there is only 1 way if jackie gets the role

jimthompson5910 (jim_thompson5910):

like I said ` We're still under the assumption George got the role. So Jackie can't get the role if George did.`

OpenStudy (pulsified333):

so what would the answer be

OpenStudy (pulsified333):

168?

jimthompson5910 (jim_thompson5910):

we're not there yet how many ways are there to fill the female role? assume that Jackie can't get the role

OpenStudy (pulsified333):

6

jimthompson5910 (jim_thompson5910):

jackie can't get the role though 6-1 = 5 so it's 5 actually

OpenStudy (pulsified333):

ah that makes sense

jimthompson5910 (jim_thompson5910):

` 5 children audition` `2 child roles available.` how many ways to fill up these slots?

OpenStudy (pulsified333):

20

jimthompson5910 (jim_thompson5910):

good

OpenStudy (pulsified333):

168*5*20 = 16800

jimthompson5910 (jim_thompson5910):

summary so far IF george gets a role, then there are... 168 ways to fill up the male roles 5 ways to fill up the female role 20 ways to fill up the child roles

jimthompson5910 (jim_thompson5910):

yes you beat me to it, you multiply those values out

jimthompson5910 (jim_thompson5910):

that's just half the answer though

OpenStudy (pulsified333):

oh

jimthompson5910 (jim_thompson5910):

we haven't considered the scenario that Jackie gets the role

jimthompson5910 (jim_thompson5910):

give it a shot and tell me what you get

OpenStudy (pulsified333):

10080 if jackie gets the role

OpenStudy (pulsified333):

do we subtract them or add them together?

jimthompson5910 (jim_thompson5910):

once again, if either Jackie or George gets the role, then the other person can't get it so we're now assuming Jackie gets the role. So george did NOT get the role. Try again

jimthompson5910 (jim_thompson5910):

if george did NOT get the role, then the pool of 9 males drops to 8

OpenStudy (pulsified333):

sorry 6720

jimthompson5910 (jim_thompson5910):

better

jimthompson5910 (jim_thompson5910):

we have 2 cases George gets the role: 16800 different ways Jackie gets the role: 6720 different ways add them up

OpenStudy (pulsified333):

okay so that is 23520

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

which is the final answer. It's the number of ways to pick all of the cast based on those specific conditions

OpenStudy (pulsified333):

it was the correct answer

OpenStudy (pulsified333):

:D

OpenStudy (pulsified333):

thanks man

jimthompson5910 (jim_thompson5910):

ok so it was a lucky guess on our part to assume order mattered

jimthompson5910 (jim_thompson5910):

no problem